PS-MPC #10 - Commissioning (Status of patrols, DSO tests, ISTs / TT2 FGC tests)

→ Europe/Zurich
Online

Online

https://cern.zoom.us/j/94981883670?pwd=UzJiZkE3TGhPV1VSTTJzZnNGK2RMUT09
Alexandre Lasheen (CERN), Denis Gerard Cotte (CERN), Frank Tecker (CERN)
    • 10:30 → 10:35
      Introduction 5m
      Speakers: Alexandre Lasheen (CERN), Denis Gerard Cotte (CERN), Frank Tecker (CERN)
    • 10:35 → 10:50
      Status of BE-ICS tests, patrols, DSO tests 15m
      Speakers: Didier Chapuis (CERN), Fernando Baltasar Dos Santos Pedrosa (CERN), Florence Pirotte (CERN), Marc Tavlet (CERN)
    • 10:50 → 11:05
      Follow-up: TT2 tests with FGC 15m
      Speaker: Denis Gerard Cotte (CERN)
    • 11:05 → 11:15
      AOB 10m

      CO readiness

      Mail from Marine, May 25, 2020 at 17:17:

      -----------------------------------

      Global status

      All services are operational now. There are a few changes in the pipeline, which are described below.  The only service whose I am waiting the status for is DIAMON/COSMOS but sure you will need them for the IST.

      -----------------------------------

      HARDWARE installations (From Benjamin):

      a/ voici ce qu’il reste avant les dates des IST :

      PSB : 

      • cablage BIX.WINJ pour les BPM
      • cablage BEX.SEJ pour les BCT d’exctraction
      • déplacer les cables des timing pour les BSW et les QSTRIPs pour qu’ils aient les bons starts
      • validation installation OASIS + Timing des DIAMON BLM
      • Deplacer cfc-400-rl4hst et l’installer au PSB cfc-361-reth6 (Magnus s’en charge)

       

      PS :

      • Installation 1 repetiteur au 269
      • Installation 1 repetiteur au 362

       

      b/ ensuite les travaux prévus pour plus tard :

      PSB :

      • OASIS CBAOS13 consolidation (attente confirmation RF)

       

      PS :

      • Cablage 5 repetiteurs au 151 (attente confirmation RF)
      • Renovation repetiteurs TOF bat 380
      • Renovation repetiteurs APRON bat 157

       

      LEIR/LN3 :

      • Cablage timing pour BCT LN3 et LEIR (attente tirage cable EN/EL)

       

      -----------------------------------

      CMW (from Wojtek): 

      • We plan to deploy a new version of the CMW Directory Service and stop all CMW Proxies next week, Wednesday 27th May , announced weeks ago. This will stop all CMW Proxies and we will block all the RDA2 communication. 
      • We don’t plan any other major changes in the following weeks

      -----------------------------------

      OASiS (from Anastasiya): 

      • Oasis will be available for the ISTs as requested.
      • This week (W22) with David we are going to deploy a new version of the front-end software relying on the dynamic PCI crate configuration. The FECs will be stable from 8 June on.
      • At the end of June - beginning of July (to be finalised and planned) we are going to deploy some bug fixes in the application server and the Oasis viewer application. The actual deployment takes few minutes and the short service interruption will be managed with the connected clients at that particular moment.

      -----------------------------------

      TIMING (from Greg): 

      • We have 2 minor bug fixes to be deployed in the LIC CT. This could be done either this or next week (to be confirmed with OP).From APIs point of view we don’t plan to change anything for the time being. 
      • We’ll likely need to deploy a new version again in a few weeks with improvements for the diagnostics (external conditions, beam evaluation), but this won’t have any impact on the system behaviour or on app developments.